A 5.00 L of a gas is collected at 100. K and then allowed to 

expand to 20.0 L. What must the new temperature be in order 

to maintain the same pressure (as required by Charles' Law)? 


1
Expert's answer
2021-10-28T02:56:31-0400

P1V1= P2V2

5×100= 20×x

500= 20X

X= 500/20

X= 25K